OpenStudy (anonymous):

MATH HELP PLEASE!!!!!!!! 1. Jose is paying for a ten dollar meal using bills in his wallet. He has five one dollar bills and four five dollar bills. If he selects two bills at random, one at a time from his wallet, without replacement, what is the probability that he will choose two five dollar bills to pay for the meal? Show your work.

jimthompson5910 (jim_thompson5910):

what is the probability he will select one 5 dollar bill?

OpenStudy (anonymous):

5/10

jimthompson5910 (jim_thompson5910):

5 one dollar bills + 4 five dollar bills ----------------- 9 bills total

jimthompson5910 (jim_thompson5910):

there are 4 things that we want out of 9 things total so the probability is actually 4/9

OpenStudy (anonymous):

is it like this? P1=5/10 P2=4/9 P(1&2)=5/10 * 4/9 = 20/90 = 2/9

jimthompson5910 (jim_thompson5910):

no

OpenStudy (anonymous):

i'm confused

jimthompson5910 (jim_thompson5910):

the probability of selecting one five dollar bill is 4/9 once you've picked a 5 dollar bill, you have 3 of them left over (out of 8 bills total now) so the probability of selecting a second 5 dollar bill is 3/8 multiply them to get: (4/9)*(3/8) = 12/72 = 1/6

jimthompson5910 (jim_thompson5910):

this means that the probability of selecting two 5 dollar bills (without replacement) is 1/6

OpenStudy (anonymous):

okay thanks

jimthompson5910 (jim_thompson5910):

yw

OpenStudy (anonymous):

I didn't really get how to do dependent events.